A business employs 125 staff, each of whom has a number from 1 to 125 assigned to them. Three are chosen to participate in a survey. A calculator generates the following three random numbers:
0.051
0.847
0.219
Find the employee numbers of the people who will be chosen to participate in the survey.

Answer by jim_thompson5910(35256)

We can multiply each of the random numbers by 125

0.051*125 = 6.375 which rounds to 6
0.847*125 = 105.875 which rounds to 106
0.219*125 = 27.375 which rounds to 27

So employees with ID numbers 6, 206 and 27 are chosen.